OpenStudy (anonymous):

what is the amplitude of h(t)=-3+2sinθ?

OpenStudy (ranga):

In y = Asin(Bx), A is the amplitude. The sine function will oscillate about the horizontal line, y = 0 (or the x-axis). In y = Asin(Bx) + k, A is still the amplitude. The sine function will oscillate about the horizontal line, y = k.

OpenStudy (anonymous):

so the amplitude is 2?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

do you also know how to find the period, vertical displacement, and horizontal displacement?

OpenStudy (ranga):

If y = Asin(Bx+C) + D, Amplitude is A. Period is (2*pi) / B Vertical displacement is D Horizontal displacement is -C/B